OpenStudy (anonymous):

Please help Verify each equation as an identity Sin^2 x - sin^2 y = cos^2 y - cos^2 x

OpenStudy (jdoe0001):

if you were to solve, say \(\bf sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)=?\) what would you get for that?